A rectangular fence is to be made by using all 64 feet of fencing. The length is 6 feet longer than the width of the fence.
Answer by fractalier(6550)

If we call the width w, then the length is w+6.
The perimeter of a rectangle is twice the length plus twice the width, so we have
(w+6) + (w+6) + w + w = 64
4w + 12 = 64
4w = 52
w = 13
w+6 = 19